OpenStudy (anonymous):

Let B be the solid whose base is the circle x^2+y^2=3^2 and whose vertical cross sections perpendicular to the x-axis are equilateral triangles. Compute the volume of B.

OpenStudy (anonymous):

The area of an equilateral triangle of side s is \[ \frac{s^2 \sqrt 3}4 \] We are going to use this in the proof http://www.mathwords.com/a/area_equilateral_triangle.htm

OpenStudy (anonymous):

The cross sections at x will be equilateral triangles of sides 2 y

OpenStudy (anonymous):

The area of each cross section is \[ \frac{ (2 y)^2 \sqrt 3}4= \sqrt 3\, y^2=\sqrt 3\, \left (9-x^2 \right ) \]

OpenStudy (anonymous):

So \[ dV=\sqrt 3\, \left (9-x^2 \right ) dx \]

OpenStudy (anonymous):

So \[ V=\int_{-3}^3 \sqrt 3\, \left (9-x^2 \right ) dx=36 \sqrt 3 \]

OpenStudy (anonymous):

what is the left and right?

OpenStudy (anonymous):

oh wait, i got it, thank you

OpenStudy (anonymous):

YW